# Self-Improving Skill Learning for Robust Skill-based Meta-Reinforcement Learning

### v3

@@ 5.2 Environmental Setup
5.3 Performance Comparison
We compare the proposed SISL with various baseline algorithms. Non-meta RL algorithms are trained directly on each test task for 0.5K iterations due to the absence of a meta-train phase. Meta-RL algorithms undergo meta-training for 10K iterations in Kitchen and Office, and 4K in Maze2D and AntMaze, followed by fine-tuning on test tasks for an additional 0.5K iterations. For SISL, the skill refinement interval $K_{\text{iter}}$ is set to 2K for Kitchen, Office, and AntMaze; 1K for Maze2D. To ensure a fair comparison, SISL counts each update process from its skill-improvement policy and high-level policy as one iteration. Table 1 presents the final average return across test tasks after the specified test iterations. The corresponding learning curves are provided in Appendix D.1 for a more detailed comparison. From the result, SAC and PEARL baselines, which do not utilize skills or offline datasets, perform poorly on long-horizon tasks, yielding a single result across all noise levels. In contrast, SPiRL, SiMPL, and SISL, which leverage skills, achieve better performance.
SPiRL and SiMPL, however, show sharp performance declines as dataset noise increases. While both perform well with expert data, SiMPL struggles under noisy conditions due to instability in its task encoder $q_{e}$, sometimes performing worse than SPiRL. Here, the baseline results for Maze2D (Expert) are somewhat lower than those reported in the SiMPL paper. This discrepancy likely arises because, in constructing the offline dataset, we considered fewer tasks compared to SiMPL, resulting in trajectories that do not fully cover the map. Interestingly, minor noise occasionally boosts performance by introducing diverse trajectories that improve skill learning, a deteailed analysis of changes in state coverage is provided in Appendix D.1. In contrast, SISL demonstrates superior robustness across all evaluated environments, consistently outperforming baselines at varying noise levels. For example, in the Kitchen environment, SISL maintains strong performance under significant noise by effectively refining useful skills, while in Maze2D, higher noise levels lead to the improvement of diverse skills, achieving perfect task completion when $\sigma=1.5$. These results highlight SISL’s ability to discover improved behavior and refine robust skills, significantly enhancing meta-RL performance. Moreover, SISL excels with both noisy and expert data, achieving superior test performance by learning more effective skills.
Fig. 6 shows the learning progress during the meta-train/test phases for Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), highlighting the performance gap between SISL and other methods. The periodic drops in SISL’s high-level performance correspond to the reinitialization of $\pi_{h}$ every $K_{\text{iter}}$. Non-meta RL algorithms, including those with RND-based exploration, struggle with long-horizon tasks, while SPiRL and SiMPL show limited improvement due to their reliance on noisy offline datasets. In contrast, SISL’s self-improving skill refinement supports continuous skill improvement, resulting in superior meta-test performance. To further assess SISL’s robustness, we perform experiments with limited offline data, random noise injection, and diverse sub-optimal datasets in Appendix D, reflecting real-world challenges such as costly data collection and unstructured anomalies. Even under these conditions, SISL consistently outperforms the baselines, demonstrating strong robustness. Furthermore, we provide a computational cost analysis with SiMPL in Appendix E, and the result shows that SISL requires about 16% more computation time per iteration during meta-training, while the meta-test phase remains unchanged. Notably, even with extended training, SiMPL fails to achieve further performance gains, highlighting SISL’s advantage.
@@ 5.3 Performance Comparison
5.4 In-depth Analysis of the Proposed Skill Refinement Process
@@
As a result, by iteration 5K in Kitchen, the agent learns to open the microwave, refining this skill to complete the task more efficiently by iteration 10K. In Maze2D, the agent explores more diverse trajectories over iterations, ultimately solving all training tasks by iteration 4K. These results highlight how SISL refines skills iteratively by leveraging prioritized data from offline and online buffers. As shown in Table 1, the learned skills generalize effectively to unseen test tasks, demonstrating SISL’s robustness and efficacy. While variations in return scales across tasks can introduce bias in reward model training, our experimental environments adopt a simple reward structure based on subtask completion, under which the reward model remains stable, as shown in Appendix F.4. For environments with more complex reward functions, per-task reward standardization can be considered. We provide $\beta$ trends and visualize refined skills, task-representation improvements, and policy skill composition in Appendix F. These analyses provide insights into SISL’s effectiveness in optimizing skill execution and enhancing task representation.
@@ 5.5 Ablation Studies
We evaluate the impact of SISL’s components and key hyperparameters in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), focusing on the effect of the prioritization temperature $T$. Additional analyses, including component evaluation and KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ for all noise levels, are in Appendix G.
Component Evaluation: To evaluate the importance of SISL’s components, we compare the meta-test performance of SISL with all components included against the following variations: (1) Without $\mathcal{B}_{\text{off}}$, relying solely on $\mathcal{B}_{\text{on}}^{i}$, to evaluate the influence of offline data in skill refinement; (2) Without $P_{\mathcal{B}_{\text{off}}}$, applying uniform sampling in $\mathcal{B}_{\text{off}}$ instead of maximum return relabeling, to evaluate the effect of prioritization; (3) Without $\mathcal{B}_{\text{on}}$, using only $\mathcal{B}_{\text{off}}$, to evaluate the contribution of high-quality samples obtained by $\pi_{\text{imp}}$ and $\pi_{h}$; and (4) Without $\pi_{\text{imp}}$, removing the skill-improvement policy, to evaluate whether exploration near the skill distribution discovers improved behaviors. Fig. 9 presents the comparison results, demonstrating significant performance drops when either buffer is removed, highlighting their critical role in effective skill discovery. Uniform sampling in $\mathcal{B}_{\text{off}}$ also reduces performance, underlining the importance of maximum return relabeling. Lastly, excluding $\pi_{\text{imp}}$ notably degrades performance, emphasizing the critical role of discovering improved behavior.
@@ 7 Conclusion
In this paper, we propose SISL, a robust skill-based meta-RL framework designed to address noisy offline demonstrations in long-horizon tasks. Through self-improving skill refinement and prioritization via maximum return relabeling, SISL effectively prioritizes task-relevant trajectories for skill learning and enabling efficient exploration and targeted skill optimization. Experimental results highlight its robustness to noise and superior performance across various environments, demonstrating its potential for scalable meta-RL in real-world applications where data quality is critical.
@@ B.3 Meta-Test Phase
After completing the meta-train phase of SISL, the meta-test phase is performed on the test task set $\mathcal{M}_{\mathrm{test}}$. In this phase, previously learned components, including the task encoder $q_{e,\theta}$, low-level skill policy $\bar{\pi}_{l,\phi}$, and skill prior $\bar{p}_{\phi}$, are kept fixed and are no longer updated. Only the high-level policy $\pi_{h,\theta}$ and high-level value function $Q_{h,\theta}$ are trained for each test task using the soft actor-critic (SAC) framework.
During meta-testing, for each test task $\mathcal{T}$, the task representation $e^{\mathcal{T}}$ is inferred from the fixed task encoder $q_{e,\theta}$. The SAC algorithm is then applied to optimize the high-level policy and value function for the specific test task, following the same loss functions as defined in Eq. (B.4) from the meta-training phase. This approach ensures efficient adaptation to unseen tasks by leveraging the fixed, pre-trained low-level skills and task representations.
@@
12      (Decoupled Policy Learning)
      Update $\pi_{\text{imp},\psi}$, $Q_{\text{imp},\psi}$ using Eq. (B.3) with $\psi\leftarrow\psi-\lambda^{\text{lr}}_{\text{imp}}\cdot\nabla_{\psi}(\mathcal{L}^{\text{critic}}_{\text{imp}}(\psi)+\mathcal{L}^{\text{actor}}_{\text{imp}}(\psi))$.
      Update $\pi_{h,\theta}$, $Q_{h,\theta}$, $q_{e,\theta}$ using Eq. (B.4) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$.
13      (Self-Improvement Skill Learning)
@@
      Update $\pi_{l,\phi}$, $q_{\phi}$, $p_{\phi}$ using Eq. (B.6) with $\phi\leftarrow\phi-\lambda^{\text{lr}}_{l}\cdot\nabla_{\phi}\mathcal{L}_{\text{skill}}(\phi)$.
15
@@
Require: Target task $\mathcal{T}$, high-level policy $\pi_{h,\theta}$, value function $Q_{h,\theta}$, task encoder $q_{e,\theta}$, and low-level policy $\pi_{l,\phi}$.
2 Collect context $c^{\mathcal{T}}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e\sim\mathcal{N}(\mathrm{0},\mathrm{I})$.
3 Compute task representation $e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})$.
@@
5   Collect high-level trajectory $\tau^{\mathcal{T}}_{h}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e^{\mathcal{T}}$.
@@ C.4 Hyperparameter Setup
For implementing the skill models ($\pi_{l},q,p$), we follow SPiRL by utilizing LSTM (Graves & Graves, 2012) for the skill encoder and MLP structures for the low-level skill policy and skill prior. For implementing the high-level models ($\pi_{h},Q_{h},q_{e}$), we follow SiMPL by utilizing Set Transformer (Lee et al., 2019) for the task encoder and MLP structures for the high-level policy and value function. Additionally, for implementing the SISL, we utilize MLP structures for $\pi_{\text{imp}}$, $Q_{\text{imp}}$, and $\hat{R}$. The detailed hidden network sizes are presented in Table C.3 and Table C.4. Table C.3 presents the network architectures (the number of nodes in fully connected layers) and the hyperparameters shared across all environments, while Table C.4 details the environment-specific hyperparameter setups.
@@
Shared
Hyperparameters
Group Name Environments
@@
High-level
Discount Factor $\gamma_{h}$ 0.99
@@
Low-level
Buffer size $\mathcal{B}^{i}_{\text{on}}$ 10K
@@
Skill-Improvement
RND state dropout ratio 0.7
@@
Environmental
Hyperparameters
Group Name Environments
@@
High-level
Buffer size $\mathcal{B}^{i}_{h}$ 3000 3000 20000 20000
@@
Low-level
Skill refinement $K_{\text{iter}}$ 2000 2000 1000 2000
@@
Skill-Improvement
Buffer size $\mathcal{B}^{i}_{\text{imp}}$ 100K 200K 100K 300K
@@
KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$
0.005 (Expert)
0.001
0.001
0.001
@@ D.2 Limited Offline Dataset Size
To further investigate the robustness of SISL, we conducted additional experiments to assess the impact of dataset size on skill learning and downstream performance. While SISL primarily addresses the challenge of refining corrupted skills from noisy demonstrations through online interaction, the size of the offline dataset remain important factors, especially in practical scenarios. To assess the impact of dataset size, we conducted additional experiments on the Kitchen environment using the same expert dataset but reduced to 50% (0.5M), 25% (0.25M), and 10% (0.1M) of its original dataset size (1M). As shown in Table D.1, SISL consistently outperforms baselines even under limited expert data, and its performance degrades more gracefully compared to baselines. These results confirm SISL’s ability to refine skills even from limited data, which is particularly valuable in real-world scenarios where collecting high-quality demonstrations is costly or infeasible.

### v4

@@ 5.3 Performance Comparison
Fig. 6 shows the learning progress during the meta-train/test phases for Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), highlighting the performance gap between SISL and other methods. The periodic drops in SISL’s high-level performance correspond to the reinitialization of $\pi_{h}$ every $K_{\text{iter}}$. Non-meta RL algorithms, including those with RND-based exploration, struggle with long-horizon tasks, while SPiRL and SiMPL show limited improvement due to their reliance on noisy offline datasets. In contrast, SISL’s self-improving skill refinement supports continuous skill improvement, resulting in superior meta-test performance. To further assess SISL’s robustness, we perform experiments with limited offline data, random noise injection, and diverse sub-optimal datasets in Appendix D, reflecting real-world challenges such as costly data collection and unstructured anomalies. Even under these conditions, SISL consistently outperforms the baselines, demonstrating strong robustness. Furthermore, we provide a computational complexity comparison of SISL, its ablation variants, and baseline methods in Appendix E. The results show that SISL requires only about 16% more computation time per iteration during meta-training, while the meta-test cost remains unchanged compared to SiMPL. Although SISL introduces the improvement policy $\pi_{\mathrm{imp}}$, the overall training cost remains similar because the total number of samples and updates matches SiMPL by splitting training tasks evenly between $\pi_{\mathrm{imp}}$ and $\pi_{h}$ for fair comparison, as described in Section 5.1. The remaining 16% overhead comes primarily from the skill refinement and reward model training. Notably, even with extended training, SiMPL fails to achieve further performance gains, highlighting SISL’s advantage.
5.4 In-depth Analysis of the Proposed Skill Refinement Process
@@
As a result, by iteration 5K in Kitchen, the agent learns to open the microwave, refining this skill to complete the task more efficiently by iteration 10K. In Maze2D, the agent explores more diverse trajectories over iterations, ultimately solving all training tasks by iteration 4K. These results highlight how SISL refines skills iteratively by leveraging prioritized data from offline and online buffers. As shown in Table 1, the learned skills generalize effectively to unseen test tasks, demonstrating SISL’s robustness and efficacy. While variations in return scales across tasks can introduce bias in reward model training, our experimental environments adopt a simple reward structure based on subtask completion, under which the reward model remains stable, as shown in Appendix F.4. For environments with more complex reward functions, per-task reward standardization can be considered. In skill-based settings, both the offline trajectories and the online training tasks share similar underlying subtasks, such as region-to-region transitions in Ant/Point environments or object-centric subtasks in Kitchen/Office domains. Because these subtasks define reward semantics consistently across training tasks and offline dataset, this relabeling process does not introduce additional instability. In addition, our softmax-based prioritization forms a distribution rather than relying on a single trajectory, which further mitigates the impact of any minor estimation error. We provide $\beta$ trends, compare refined skills via skill trajectory visualization, task-representation improvements, and policy skill composition in Appendix F. These analyses provide insights into SISL’s effectiveness in optimizing skill execution and enhancing task representation.
@@ 5.5 Ablation Studies
We evaluate the impact of SISL’s components and key hyperparameters in Kitchen ($\sigma=0.3$) and Maze2D ($\sigma=1.5$), focusing on the effect of the prioritization temperature $T$.
Component Evaluation: To evaluate the importance of SISL’s components, we compare the meta-test performance of SISL with all components included against the following variations: (1) Without $\mathcal{B}_{\text{off}}$, relying solely on $\mathcal{B}_{\text{on}}^{i}$, assessing the influence of offline data in skill refinement; (2) Without $P_{\mathcal{B}_{\text{off}}}$, applying uniform sampling in $\mathcal{B}_{\text{off}}$ instead of maximum return relabeling, to test the effect of prioritization; (3) Without $\mathcal{B}_{\text{on}}$, using only $\mathcal{B}_{\text{off}}$, examining the contribution of high-quality samples obtained by $\pi_{\text{imp}}$ and $\pi_{h}$; and (4) Without $\pi_{\text{imp}}$, removing the skill-improvement policy, verifying whether exploration near the skill distribution discovers improved behaviors. Fig. 9 presents the comparison results, demonstrating significant performance drops when either buffer is removed, highlighting their critical role in effective skill discovery. Uniform sampling in $\mathcal{B}_{\text{off}}$ also reduces performance, underlining the importance of maximum return relabeling. Lastly, excluding $\pi_{\text{imp}}$ notably degrades performance, emphasizing the critical role of discovering improved behavior.
To provide a clearer understanding of SISL’s components, we include several additional ablations in the Appendix. First, we analyze the temperature $T$ for all environments and noise levels in Appendix G.2, showing that its optimal value depends mainly on the return scale of each environment. Second, we study the KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ in Appendix G.3 and find that performance is stable across a wide range of values, but drops sharply when the coefficient is zero because the guiding effect of the KL term disappears. Third, we examine the reinitialization interval $K_{\text{iter}}$ in Appendix G.5, confirming that the high-level policy needs a minimum amount of adaptation time after each refinement step, while larger intervals produce similar outcomes as long as skills are updated a few times during training. Together, these results highlight the role and influence of each component in SISL.
@@ 7 Conclusion
In this paper, we propose SISL, a robust skill-based meta-RL framework designed to address noisy offline demonstrations in long-horizon tasks. Through self-improving skill refinement and prioritization via maximum return relabeling, SISL effectively prioritizes task-relevant trajectories for skill learning and enables efficient exploration and targeted skill optimization. Experimental results highlight its robustness to noise and superior performance across various environments, demonstrating its potential for scalable meta-RL in real-world applications where data quality is critical.
Acknowledgment
This work was supported partly by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220469, Development of Core Technologies for Task-oriented Reinforcement Learning for Commercialization of Autonomous Drones), (No. RS-2025-25442824, AI Star-Fellowship Program (UNIST)), and (No. RS-2020-II201336, Artificial Intelligence Graduate School Support (UNIST)), and partly by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-23523191, LLM-Based Multi-Agent Reinforcement Learning for End-to-End Large Autonomous Swarm Control).
@@ B.2 Meta-Train Phase
After the meta-train phase is completed, the final meta-train phase parameter is stored as $\theta_{\text{final}}\leftarrow\theta$ and is subsequently used during the meta-test phase.
@@ B.3 Meta-Test Phase
After completing the meta-train phase of SISL, the meta-test phase is performed on the test task set $\mathcal{M}_{\mathrm{test}}$. In this phase, previously learned components, including the low-level skill policy $\bar{\pi}_{l,\phi}$, skill prior $\bar{p}_{\phi}$, and task encoder $q_{e,\theta_{\text{final}}}$, are kept fixed and are no longer updated. Only the high-level policy $\pi_{h,\theta}$ and high-level value function $Q_{h,\theta}$ are trained for each test task using the soft actor-critic (SAC) framework.
During meta-testing, for each test task $\mathcal{T}$, the task representation $e^{\mathcal{T}}$ is inferred from the fixed task encoder $q_{e,\theta_{\text{final}}}$. The SAC algorithm is then applied to optimize the high-level policy and value function for the specific test task, following the same loss functions as defined in Eq. (B.4) from the meta-training phase. This approach ensures efficient adaptation to unseen tasks by leveraging the fixed, pre-trained low-level skills and task representations.
@@
(Initial Skill Learning)
Update $\pi_{l,\phi}$, $q_{\phi}$, $p_{\phi}$ using Eq. (B.1) with $\phi\leftarrow\phi-\lambda^{\text{lr}}_{l}\cdot\nabla_{\phi}\mathcal{L}_{\text{spirl}}(\phi)$.
@@
      (Decoupled Policy Learning)
12      Update $\pi_{\text{imp},\psi}$, $Q_{\text{imp},\psi}$ using Eq. (B.3) with $\psi\leftarrow\psi-\lambda^{\text{lr}}_{\text{imp}}\cdot\nabla_{\psi}(\mathcal{L}^{\text{critic}}_{\text{imp}}(\psi)+\mathcal{L}^{\text{actor}}_{\text{imp}}(\psi))$.
13      Update $\pi_{h,\theta}$, $Q_{h,\theta}$, $q_{e,\theta}$ using Eq. (B.4) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$.
      (Self-Improvement Skill Learning)
@@
15      Update $\pi_{l,\phi}$, $q_{\phi}$, $p_{\phi}$ using Eq. (B.6) with $\phi\leftarrow\phi-\lambda^{\text{lr}}_{l}\cdot\nabla_{\phi}\mathcal{L}_{\text{skill}}(\phi)$.
16
@@
21
Save the final meta-train phase parameter $\theta_{\text{final}}\leftarrow\theta$.
@@
Require: Target task $\mathcal{T}$, high-level policy $\pi_{h,\theta}$, value function $Q_{h,\theta}$, task encoder $q_{e,\theta_{\text{final}}}$, low-level policy $\bar{\pi}_{l,\phi}$ , and skill prior $\bar{p}_{\phi}$.
2 Collect context $c^{\mathcal{T}}$ using $\pi_{h,\theta}$ with $\bar{\pi}_{l,\phi}$, $e\sim\mathcal{N}(\mathrm{0},\mathrm{I})$.
3 Compute task representation $e^{\mathcal{T}}\sim q_{e,\theta_{\text{final}}}(\cdot|c^{\mathcal{T}})$.
@@
5   Collect high-level trajectory $\tau^{\mathcal{T}}_{h}$ using $\pi_{h,\theta}$ with $\bar{\pi}_{l,\phi}$, $e^{\mathcal{T}}$.
@@ C.4 Hyperparameter Setup
For implementing the skill models ($\pi_{l},q,p$), we follow SPiRL by utilizing LSTM (Graves and Graves, 2012) for the skill encoder and MLP structures for the low-level skill policy and skill prior. For implementing the high-level models ($\pi_{h},Q_{h},q_{e}$), we follow SiMPL by utilizing Set Transformer (Lee et al., 2019) for the task encoder and MLP structures for the high-level policy and value function. Additionally, for implementing the SISL, we utilize MLP structures for $\pi_{\text{imp}}$, $Q_{\text{imp}}$, and $\hat{R}$. The detailed hidden network sizes are presented in Table C.3 and Table C.4. Table C.3 presents the network architectures (the number of nodes in fully connected layers) and the hyperparameters shared across all environments, while Table C.4 details the environment-specific hyperparameter setups.
@@
Shared Hyperparameters Group Name Environments
@@
High-level Discount Factor $\gamma_{h}$ 0.99
@@
Low-level Buffer size $\mathcal{B}^{i}_{\text{on}}$ 10K
@@
Skill-Improvement RND state dropout ratio 0.7
@@
Environmental Hyperparameters Group Name Environments
@@
High-level Buffer size $\mathcal{B}^{i}_{h}$ 3000 3000 20000 20000
@@
Low-level Skill refinement $K_{\text{iter}}$ 2000 2000 1000 2000
@@
Skill-Improvement Buffer size $\mathcal{B}^{i}_{\text{imp}}$ 100K 200K 100K 300K
@@
KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ 0.005 (Expert) 0.001 0.001 0.001
@@ D.2 Limited Offline Dataset Size
To further investigate the robustness of SISL, we conducted additional experiments to assess the impact of dataset size on skill learning and downstream performance. While SISL primarily addresses the challenge of refining corrupted skills from noisy demonstrations through online interaction, the size of the offline dataset remains an important factor, especially in practical scenarios. To assess the impact of dataset size, we conducted additional experiments on the Kitchen environment using the same expert dataset but reduced to 50% (0.5M), 25% (0.25M), and 10% (0.1M) of its original dataset size (1M). As shown in Table D.1, SISL consistently outperforms baselines even under limited expert data, and its performance degrades more gracefully compared to baselines. These results confirm SISL’s ability to refine skills even from limited data, which is particularly valuable in real-world scenarios where collecting high-quality demonstrations is costly or infeasible.